NOAA-15 , HIRS, HIRS Anomaly Update, Implementation Date 17 April 2005 - Issued: 18 April 2005, 1745 UTC

Subject: NOAA-15 , HIRS, HIRS Anomaly Update, Implementation Date 17 April 2005 - Issued: 18 April 2005, 1745 UTC
"Emily Harrod"
Mon, 18 Apr 2005 13:45:46 -0400

NOAA-15 HIRS Filter Wheel Problem Continues

Date/Time(UTC) Message Issued:
18 April 2005/1745 UTC

Satellite(s) Involved:

Instrument(s) Involved:

Product(s) Involved/Affected:
All Products using HIRS instrument on NOAA-15

Date/Time(UTC) of Initial Implementation:
17 January 2005, 1500 UTC

Details/Specifics of Change:
Official notification from SOCC reports: "At approximately 13:00Z on Sunday, 17 Apr, the N15 HIRS filter motor current soared to red high limits of 320+ mAmps. This surge was accompanied by greatly increased temperatures as well as very erratic period monitor values and intermittent loss of filter wheel synch. At 14:52Z, ..., EMOSS engineers turned on the HIRS filter motor heater. Shortly thereafter, the filter motor current surged again to its highest levels ever (408 mAmps) with accompanying high temperatures and a steady loss of filter wheel synch. Plans are to turn the heater back off on Monday, 14:20Z to determine if this will improve the HIRS situation. If it does not, a downgrade in status to RED is under consideration. Additional information will be provided as it is obtained. It is assumed that all HIRS products are degraded at this time..."

Follow-up report: "Imagery was OK at the start of the major surge (17 Apr 13:00Z), and completely corrupted on all channels on the next pass at 14:40Z (the same pass heater turnon occurred) and remained corrupted through the most recent pass today at 14:20Z after heater was turned back off."

SOCC engineers continue to monitor and investigate the anomaly.

Contact Person(s) Name/Email/Phone Number for Questions:
Renee Smith-Dearing, 301-817-4117, - SOCC POES Engineering
Emily Harrod, 301-452-5247 ext 117, - IPD operations support

Web site(s) that contain other relevant information (where applicable): - History of level 1B notices - All user notices - SOCC polar satellites status - SOCC morning reports

"The contents of this message are mine personally and do not necessarily reflect any position of the Government or the National Oceanic and Atmospheric Administration."

* Emily D. Harrod
* Computer Operations Branch
* Pre-Product Processing Group
* E/SP11, Federal Building #4, room 0318
* 5200 Auth Road
* Suitland, Maryland 20746-4304
* voice 301-457-5247 ext 117
* fax 301-457-5199
* email